OpenStudy (anonymous):

calc help, Derivative of G(x) = sec^2 (xe^(x^(2)))

OpenStudy (anonymous):

Put \[ u = x e^{x^2} \]

OpenStudy (anonymous):

yes d/dx(\sec^2(e^(x^2) x)) Use the chain rule, d/dx(\sec^2(e^(x^2) x)) = ( du^2)/( du) ( du)/( dx), where u = \sec(e^(x^2) x) and ( du^2)/( du) = 2 u: = | 2 \sec(e^(x^2) x) (d/dx(\sec(e^(x^2) x))) | Use the chain rule, d/dx(\sec(e^(x^2) x)) = ( dsec(u))/( du) ( du)/( dx), where u = e^(x^2) x and ( dsec(u))/( du) = \sec(u) \tan(u): = | 2 \sec(e^(x^2) x) (\tan(e^(x^2) x) \sec(e^(x^2) x) (d/dx(e^(x^2) x))) | Use the product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = e^(x^2) and v = x: = | 2 \tan(e^(x^2) x) \sec^2(e^(x^2) x) (e^(x^2) (d/dx(x))+x (d/dx(e^(x^2)))) | Use the chain rule, d/dx(e^(x^2)) = ( de^u)/( du) ( du)/( dx), where u = x^2 and ( de^u)/( du) = e^u:= | 2 \tan(e^(x^2) x) \sec^2(e^(x^2) x) (x (e^(x^2) (d/dx(x^2)))+e^(x^2) (d/dx(x))) The derivative of x is 1 = 2 \tan(e^(x^2) x) \sec^2(e^(x^2) x) (e^(x^2) x (d/dx(x^2))+e^(x^2)) | The derivative of x^2 is 2 x: = | 2 (e^(x^2) x (2 x)+e^(x^2)) \tan(e^(x^2) x) \sec^2(e^(x^2) x)\] FINALLY d/dx(sec^2(x e^(x^2))) = 2 e^(x^2) (2 x^2+1) tan(e^(x^2) x) sec^2(e^(x^2) x)

OpenStudy (anonymous):

Power rule for Differentiation: for: \[f(x) = x^n\] then: \[f'(x)=\frac{d}{dx} n * x^{n-1)} \] Basically bring the exponent down as a coefficient, then subtract the exponent by one. Chain Rule: for: \[u^n \] then: \[(n*u^{n-1})*\frac{d}{du}u\] Do the power rule, then take the derivative of what's inside the "shell"/"layer"/"whatever"

OpenStudy (anonymous):

sec^2 (xe^(x^(2))) \[ G'[x] = 2 \sec (u) (\sec(x) \tan(x) )u'(x) \]

OpenStudy (anonymous):

writen so far 2tan(xe^(x^(2)))sec^2(xe^(x^(2))) * 1e^(x^(2))+x * ?

OpenStudy (anonymous):

\[ G'[x] = 2 \sec (u) (\sec(x) \tan(x) )u'(x)=\\ 2 \sec (x e^{x^2}) (\sec(x) \tan(x) )( e^{x^2} + 2 x^2 e^{x^2}) \]